# Euclid Quick Data Release (Q1) - Spectroscopic search, classification and analysis of ultracool dwarfs in the Deep Fields

### v1

@@ Abstract
The Near-Infrared Spectrometer and Photometer onboard the Euclid space mission has obtained near-infrared (NIR) slitless spectra of millions of objects, including hundreds of ultracool dwarfs. Euclid observations retrieve images and spectra simultaneously. This observing mode marks a new era in the discovery of new objects, such as L- and T-type dwarfs, which can be found from direct identification through the H2O and CH4 absorption bands. NISP spectral resolution ( 450) is enough to classify the objects by the spectral type using known standard templates. Q1 provided more than 4 million NIR spectra in one visit to the Euclid Deep Fields. The large amount of spectra released in these fields allowed us to: a) confirm almost half of the UCD photometric candidates by J. Y. Zhang et al. (2024); b) discover at least 10 new late L- and T-type dwarfs searching directly in the spectra database; and c) spectroscopically confirm one hundred more candidates from a new photometric selection conducted by M. Žerjal et al. (2025). We present a preliminary list of Euclid UCD templates built by the combination of the best spectra from all these searches. We include the first spectral analysis of confirmed UCDs from Q1 data; spectral classifications; determination of effective temperatures; H2O, CH4 and NH3 spectral indices; and measurements of the KI absorption doublet. This paper is a first step in the study of Euclid UCDs and will be improved with each subsequent data release.
@@ 1 Introduction
The advent of the Euclid space mission (Y. Euclid Collaboration: Mellier et al., 2024) is expected to foster major advancements in many fields of astrophysics. The Euclid Quick Data Release (Q1) includes slitless near-infrared (NIR) spectra of 4.3 million objects observed with the Near-Infrared Spectrometer and Photometer (NISP; K. Euclid Collaboration: Jahnke et al., 2024). Among those millions of sources, several hundred are UCDs, which have effective temperatures below 2700 K (J. D. Kirkpatrick et al., 1995) and down to 250 K for the coldest object observed to date (K. L. Luhman, 2014; K. L. Luhman et al., 2024). The spectral classification system for UCDs follows their decreasing temperatures: late M-type (M7 and later, J. D. Kirkpatrick et al. 1997, L-type (E. L. Martin et al., 1997, 1998, 1999; J. D. Kirkpatrick et al., 1999, 2000), T-type (A. J. Burgasser et al., 2002; T. R. Geballe et al., 2002; A. J. Burgasser et al., 2006), and Y-type dwarfs (P. Delorme et al., 2008; M. C. Cushing et al., 2011; J. D. Kirkpatrick et al., 2011).
This paper focuses on the spectroscopic capabilities of Euclid in the field of UCD science, and is part of a collection of papers dedicated to the study of ultracool dwarfs. Here we investigate whether NISP’s spectral resolution and sensitivity are sufficient to analyze the molecular bands of H2O, CH4 and NH3 (see E. L. Martín et al. 2021), as well as other spectral features, in order to classify them and to study their physical properties. To this end, we focus on four objectives:
To confirm the photometric UCD candidates published by J. Y. Zhang et al. (2024), through the analysis of Euclid spectra.
To discover new late-L and T-type UCDs by directly searching in the spectra database.
To confirm the photometric UCD candidates selected from Q1 data by M. Žerjal et al. (2025).
@@
Q1 (H. Aussel et al., 2025) provides calibrated data of one visit to the EDFs. The data include coadded and calibrated images and extracted spectra, as well as catalogs for the different type of objects (stars, galaxies, etc.). Q1 data include 4.3 million extracted NISP spectra which cover the wavelength range 1200 to 1900 nm (see Y. Euclid Collaboration: Copin et al., 2025; V. Le Brun et al., 2025). The spectra files include wavelength (in air), flux density, the flux density variance, along with the following flags 111http://st-dm.pages.euclid-sgs.uk/data-product-doc/dm10/:
@@
While slitless spectra observing mode provides the spectrum of every object in the field of view, there is a significant probability of contamination among the objects. In addition, given that NISP has a mosaic of 16 detectors with gaps between them, all or part of the object spectrum could fall into these gaps for some of the dithering positions (out of four in a single ROS). An example of a Q1 spectrum of E597913 (alias composed by the first 6 digits of the Euclid ID) of a confirmed T7 UCD in the EDF-N identified by J. Y. Zhang et al. (2024) is shown in Fig. 1. The spectrum is displayed in two colors: green lines represent the spectra with NDITH 2 and red lines represent the spectra with QUALITY 0.7. The number of dithering positions can change with wavelength, so plotting the spectrum in two colors helps identify any unreliable portions (i.e. from unidentified contamination). We chose the value of 0.7 for QUALITY as it filters most of the artifacts. Future data releases will be based on the combination of more dithering positions, so that QUALITY and NDITH values will increase.
@@ 2 Retrieving UCD candidates from Q1 data
Two photometric catalogs and one direct search are the main input to this work:
The photometric catalog of J. Y. Zhang et al. (2024, hereafter Zhang’s compilation) is a selection of previously known objects filtered to find UCD candidates. Spectroscopically confirmed UCDs in this work are identified as “UCD benchmarks”.
The spectroscopic search of late L- and T-type objects in the 4.3 million spectra.
The photometric catalog of Žerjal et al. (in prep., hereafter Žerjal’s catalog) which has been prepared using the UCD benchmarks.
@@ 2.1 Zhang’s compilation
In preparation for Euclid observations, J. Y. Zhang et al. (2024) performed a photometric selection of UCDs in the region of the EDFs, and published a list of 515 candidates. They selected eight candidates which they confirmed as UCDs by obtaining reconnaissance spectra from the 8.2-m Very Large Telescope (VLT) and 10.4-m Gran Telescopio Canarias (GTC). Five of the objects from this selection are found in Q1 data (our alias in parentheses): WISEA J035909.75474056.8 (E597913), WISEA J035231.80491059.4 (E581332), WISEA J033234.35273333.8 (E531434), WISEA J040254.85470440.9 (E607287), and WISEA J175730.71+674138.4 (E269377).
Zhang’s compilation includes NIR photometry from the Two Micron All Sky Survey (2MASS; R. M. Cutri et al., 2003; M. F. Skrutskie et al., 2006). In this work we label these photometric measurements with the “2MASS” suffix (e.g., ) to differentiate from NISP photometry ( , , and ).
@@
2.2 Spectroscopic search for late-L and T dwarfs
Given that Zhang’s compilation had a very low number of late L- and T-type candidates, and that Q1 data released 4.3 million spectra, we decided to search for the accentuated spectral features of late-L and T dwarfs directly in the spectra database without any photometric analysis. In particular, the combined absorption of H2O and CH4 is so evident in the spectral range covered by NISP that a fast algorithm computing ratios can easily detect these absorption bands. We developed an algorithm that filters the data to eliminate unreliable values using QUALITY and NDITH flags and computes three ratios from four narrow spectral regions. The values for these regions were adapted from the NIR spectral indices defined by A. J. Burgasser et al. (2006), to the expected Euclid spectra of T dwarfs. These regions are plotted together with four T-dwarf standard templates in Fig. 2. The regions are labeled as P and P for the peak regions, and and for deepest region of the water absorption bands. We selected these regions because Euclid observations are free from telluric absorption bands. The search algorithm normalize the integrated fluxes over the regions to the peak values and computes three ratios P-, P-, and P-. It works in two steps, first with the spectra free from unreliable values, then for the spectra convolved with a gaussian profile. In both cases it filters by the minimum values of 0.6 and 0.5 for P- and P-, respectively. In the second step it filters by P- 0.35. We calibrated these values running the search algorithm on the spectra of known T dwarfs. We ran the search algorithm over the over the entire database with satisfying results. The output included numerous spectra filled with contamination from other sources that resembled portions of expected spectral energy distribution. After a visual inspection, we kept the best findings, which turned into 10 new UCDs not published before. Table 1 list these 10 objects discovered by the search algorithm, and gives the values of the search indices. This table includes their spectral type, described in Sect. 3.1. It is interesting to see that the 10 objects spectral types span homogeneously from L9 to T5. The objects retrieved by the algorithm that were not discoveries, are described in Sect. 3.3.
@@ 2.2 Spectroscopic search for late-L and T dwarfs
One of the main advantages of Euclid observations is the simultaneous retrieval of images and spectra. This observing mode marks a new era in catalog creation, because the point-source filtering criteria can be tuned based on the immediate spectroscopic confirmation (Žerjal et al. in prep.). While this confirmation can be limited by the spectrograph sensitivity or availability, the selection criteria can be extrapolated to the rest of objects, reaching the photometric limiting magnitudes. Using the UCD benchmarks and the direct-search T-type objects to verify the selection criteria, M. Žerjal et al. (2025) created a catalog with more than 5000 UCD candidates. The photometric selection could be biased and limited by the sensitivity of the VIS instrument as described in the same paper.
@@ 3.1 Spectral type (SpT) classification
We classified the objects by comparing the spectra to the standard templates compiled in SPLAT (A. J. Burgasser & Splat Development Team, 2017), and selecting the spectral type by the best fit in the full wavelength range. Appendix A gives the details of the standard templates employed. The result for a selection of the best quality spectra (few artifacts and high signal to noise, where possible) for each spectral subclass is plotted in Fig. 3 to show the spectral sequence. These spectra are plotted without any smoothing. The missing values are due to artefact removal (as shown in red in Fig. 1).
@@ 3.2 Spectra for Zhang’s compilation
We searched for Zhang’s compilation of UCD candidates in the Q1 released spectra using their object coordinates. Out of 4.3 million spectra, we found 458 objects matching within 1 arcsecond and 52 objects matching within 1 to 2.5 arcsecond of the coordinates. Together, these 510 objects account for 99% of Zhang’s compilation. An exhaustive comparison of these spectra with standard templates allowed us to confirm 224 UCDs. Table 2 shows the number of cross-matched objects, the number of confirmed dwarfs and the number of UCDs among them. In the case of EDF-N the confirmed UCDs represent 35% of the candidates with available spectra, given the limited depth of Zhang’s compilation in this field. These numbers represent a minimum, as we excluded some candidates from confirmation owing to the poor SNR of their spectra. Future visits to the EDFs will allow us to increase the total number of confirmations. Fig. 4 shows the distribution of the number of objects per magnitude. The histograms are divided by field (North, South, and Fornax) to show the objects confirmed spectroscopically in each case. Note that from Zhang’s compilation, EDF-N is the more limited field in magnitude. The UCDs with the best fits to the standard templates (negligible fit residuals) are selected as the ”UCD benchmarks”. There are 60 objects in this list, 38 late-M dwarfs, 19 L dwarfs, and 3 T dwarfs with magnitudes 14 19. Table 6 in Appendix B provides the full list.
The rest of the confirmed objects include 55 peculiar late M- and L-dwarfs, 70 with large residuals and 39 that look like L dwarfs but cannot be reliably classified at the moment. The two latter group of objects must wait for future visits to the EDFs which will allow a better classification, as the survey will gradually gain an additional depth of 2 magnitudes upon completion (Y. Euclid Collaboration: Mellier et al., 2024).
@@
Fig. 5 shows the spectra of E535151, the faintest L1 UCD from Zhang’s compilation found in Q1 data, together with that of E654087, the brightest L1 UCD from the same compilation. For this spectral type the span goes from 15.1 18.6.
3.3 Spectra retrieved by the spectroscopic search algorithm
@@
The 10 new UCDs discovered by the search algorithm were spectroscopically classified and their SpT is included in Table 1. There is one UCD for each spectral subtype, from L9 to T5. There are also two objects, E644720 and E265716, that have a peculiar spectral energy distribution (signaled with a “?” symbol), as shown in Fig. 6. E265716 could be a binary not revealed by the images, owing to its peculiar spectrum that fits to a T4 in the -band and T2 in -band. We are investigating this possibility. There are 5 additional T dwarfs found by the algorithm that were already known: one from Zhang’s compilation, two found by Žerjal’s et al. (in prep.) in the process of creating their photometric catalog, one from G. N. Mace et al. (2013), and one found as a QSO contaminant by A. Mohandasan et al. (2025). These objects are included in Table 7 (Appendix C) with the reference indicated next to the alias.

### v2

@@ Abstract
The Near-Infrared Spectrometer and Photometer onboard the Euclid space mission has obtained near-infrared (NIR) slitless spectra of millions of objects, including hundreds of ultracool dwarfs. Euclid observations retrieve images and spectra simultaneously. This observing mode marks a new era in the discovery of new objects, such as L- and T-type dwarfs, which can be found from direct identification through the H2O and CH4 absorption bands. NISP spectral resolution ( 450) is enough to classify the objects by the spectral type using known standard templates. Q1 provided more than 4 million NIR spectra in one visit to the Euclid Deep Fields. The large amount of spectra released in these fields allowed us to: a) confirm the ultracool dwarf nature of almost half of the photometric candidates compiled by J. Y. Zhang et al. (2024); b) discover at least 11 new late L- and T-type dwarfs by a specific spectral index search in Q1 data; and c) spectroscopically confirm one hundred more candidates from a new photometric selection conducted by Žerjal et al. (in prep.). We present a preliminary list of Euclid ultracool dwarf templates built by the combination of the best spectra from all these searches. We include the first spectral analysis of confirmed ultracool dwarfs from Q1 data; spectral classifications; determination of effective temperatures; H2O, CH4 and NH3 spectral indices; and measurements of the KI absorption doublet. This paper is a first step in the study of Euclid ultracool dwarfs and will be improved with each subsequent data release.
@@ 1 Introduction
The advent of the Euclid space mission (Y. Euclid Collaboration: Mellier et al., 2024) is expected to foster major advancements in many fields of astrophysics. The Euclid Quick Data Release (Q1) includes slitless near-infrared (NIR) spectra of 4.3 million objects observed with the Near-Infrared Spectrometer and Photometer (NISP; K. Euclid Collaboration: Jahnke et al., 2024). Among those millions of sources, several hundred are ultracool dwarfs (UCDs), which have effective temperatures below 2700 K (J. D. Kirkpatrick et al., 1995) and down to 250 K for the coldest object observed to date (K. L. Luhman, 2014; K. L. Luhman et al., 2024). The spectral classification system for UCDs follows their decreasing temperatures: late M-type (M7 and later, J. D. Kirkpatrick et al. 1997, L-type (E. L. Martin et al., 1997, 1998, 1999; J. D. Kirkpatrick et al., 1999, 2000), T-type (A. J. Burgasser et al., 2002; T. R. Geballe et al., 2002; A. J. Burgasser et al., 2006), and Y-type dwarfs (P. Delorme et al., 2008; M. C. Cushing et al., 2011; J. D. Kirkpatrick et al., 2011).
This paper focuses on the spectroscopic capabilities of Euclid in the field of ultracool dwarf (UCD) science, and is part of a collection of papers dedicated to the study of UCDs. Here we investigate whether NISP’s spectral resolution and sensitivity are sufficient to analyze the molecular bands of H2O, CH4 and NH3 (see E. L. Mart´ın et al. 2021), as well as other spectral features, in order to classify them and to study their physical properties. To this end, we focus on four objectives:
To confirm the ultracool dwarf nature of the photometric candidates published by J. Y. Zhang et al. (2024), through the analysis of Euclid spectra.
To discover new late-L and T-type UCDs by a specific spectral index search in the full Q1 database.
To confirm the photometric UCD candidates selected from Q1 data by Žerjal et al. (in prep.).
@@
Q1 (H. Aussel et al., 2025) provides calibrated data of one visit to the EDFs. The data include coadded and calibrated images and extracted spectra, as well as catalogs for the different type of objects (stars, galaxies, etc.). Q1 data include 4.3 million extracted NISP spectra which cover the wavelength range 1200 to 1900 nm (see Y. Euclid Collaboration: Copin et al., 2025; V. Le Brun et al., 2025). The spectra files include wavelength (in air), flux density, the flux density variance, along with the following flags 111https://euclid.esac.esa.int/dr/q1/dpdd/:
@@
While slitless spectra observing mode provides the spectrum of every object in the field of view, there is a significant probability of contamination among the objects. In addition, given that NISP has a mosaic of 16 detectors with gaps between them, all or part of the object spectrum could fall into these gaps for some of the dithering positions (out of four in a single ROS). An example of a Q1 spectrum of E597913 (alias composed by the first 6 digits of the Euclid ID) of a confirmed T7 UCD in the EDF-N identified by J. Y. Zhang et al. (2024) is shown in Fig. 1. The spectrum is displayed in two colors: green lines represent the spectra with NDITH 2 and magenta lines represent the spectra with QUALITY 0.7. The number of dithering positions can change with wavelength, so plotting the spectrum in two colors helps to identify any unreliable portions (i.e. from unidentified contamination). We chose the value of 0.7 for QUALITY as it filters most of the artifacts. Future data releases will be based on the combination of more dithering positions, so that QUALITY and NDITH values will increase.
@@ 2 Retrieving UCD candidates from Q1 data
We have selected UCD candidates from a previously published photometric catalog and from an independent spectral index search. The previous photometric catalog is taken from J. Y. Zhang et al. (2024, hereafter Zhang’s compilation). It is a selection of known objects filtered to find UCD candidates. The ones spectroscopically confirmed as UCDs by this work are called hereafter “UCD benchmarks”. Our independent spectral index search of late L- and T-type objects in the 4.3 million Q1 spectra adds some more UCDs. And, a new photometric catalog is currently in preparation using the UCD benchmarks and the feedback from this work. Its is called hereafter Žerjal’s catalog (Žerjal et al., in prep.).
@@ 2.1 Zhang’s compilation
In preparation for Euclid observations, J. Y. Zhang et al. (2024) performed a photometric selection of UCDs in the region of the EDF-N, and selected candidates from the catalogs published by C. Reylé (2018) and A. Carnero Rosell et al. (2019) for the EDF-F and EDF-S. Zhang’s compilation includes a total of 515 candidates.
They selected eight candidates which they confirmed as UCDs by obtaining reconnaissance spectra from the 8.2-m Very Large Telescope (VLT) and 10.4-m Gran Telescopio Canarias (GTC). Five of the objects from this selection are found in Q1 data (our alias in parentheses): WISEA J035909.75474056.8 (E597913), WISEA J035231.80491059.4 (E581332), WISEA J033234.35273333.8 (E531434), WISEA J040254.85470440.9 (E607287), and WISEA J175730.71+674138.4 (E269377).
Zhang’s compilation includes NIR photometry from the Two Micron All Sky Survey (2MASS; R. M. Cutri et al., 2003; M. F. Skrutskie et al., 2006) for all the objects in the EDF-N, and for a few ones in the EDF-F and EDF-S. The latter two fields have most of their photometry from the VISTA Hemisphere Survey (VHS; R. G. McMahon et al., 2013). The use of different sources produces a difference in depth between the sample for EDF-N and that for the other two fields, as we will see in Sect. 3.2.
In this work we label these photometric measurements without any suffix (e.g., ) to differentiate from NISP photometry ( , , and ).
@@
| E271934 | L9p | 0.7 | 0.6 | 0.4 |
2.2 Spectral index search for late-L and T dwarfs
Given that Zhang’s compilation had a very low number of late L- and T-type candidates, and that Q1 data released 4.3 million spectra, we decided to search for the accentuated spectral features of late-L and T dwarfs directly in the spectra database without any photometric analysis. In particular, the combined absorption of H2O and CH4 is so evident in the spectral range covered by NISP that a fast algorithm computing ratios can easily detect these absorption bands. We developed an algorithm that filters the data to eliminate unreliable values using QUALITY and NDITH flags and computes three ratios from four narrow spectral regions. The values for these regions were adapted from the NIR spectral indices defined by A. J. Burgasser et al. (2006), to the expected Euclid spectra of T dwarfs. These regions are plotted together with four T-dwarf standard templates in Fig. 2. The regions are labeled as P and P for the peak regions, and and for the deepest region of the water absorption bands. We selected these regions because Euclid observations are free from telluric absorption bands. The search algorithm normalizes the integrated fluxes over the regions to the peak values and computes three ratios P-, P-, and P-. It works in two steps, first with the spectra free from unreliable values, then for the spectra convolved with a gaussian profile (boxcar = 7). In both cases it filters by the minimum values of 0.6 and 0.5 for P- and P-, respectively. In the second step it filters by P- 0.35. We calibrated these values running the search algorithm on the spectra of known T dwarfs. We ran the search algorithm over the entire database with satisfying results. The output included numerous spectra filled with contamination from other sources that resembled portions of expected spectral energy distribution. After a visual inspection, we kept the best findings, which turned into 11 new UCDs not published before. Table 1 lists these 11 objects discovered by the search algorithm, and gives the values of the search indices. This table includes their spectral type, described in Sect. 3.1. It is interesting to see that the 11 objects spectral types span homogeneously from L9 to T5. The objects retrieved by the algorithm that were previously known, are described in Sect. 3.3. We also conducted a search for Y dwarfs, with no concluding results.
@@ 2.2 Spectral index search for late-L and T dwarfs
A new catalog of photometric UCD candidates from Q1 is under preparation by Žerjal et al. (in prep.). Special attention has been given to the selection of the point-source objects, as the vast majority of Euclid detections are extragalactic. The first step is focused on isolating point sources based on morphological criteria and relatively strict data quality cuts for each object (SNR 3 in the , , and bands). These cuts reduce the sample to just under 2% of the original Q1 dataset. UCD candidates are then selected using a color–color diagram ( - vs. - ). The inclusion of the optical band is crucial, as it allows UCDs, especially T dwarfs, to stand out from the rest of the sample. Guided by the UCD benchmarks, the selection criterion is set to - 2.5, a typical color for late-M dwarfs. To date, this selection has yielded more than 5000 UCD candidates, some of which are included in this paper.
@@ 3.1 Spectral type (SpT) classification
We classified the objects by comparing the spectra to the standard templates compiled in SPLAT (A. J. Burgasser & Splat Development Team, 2017) and selecting the spectral type by the best fit. Appendix A gives the details of the standard templates. Owing to the reduced wavelength range and the presence of artifacts, the best fit is obtained by two estimation methods. The first one is by minimizing chi-squared values over the full wavelength range. The second method is by residual minimization over four wavelength ranges defined by NISP wavelength range and the telluric absorption bands present in the standard templates: 12101350 nm, 13501530 nm, 15301800 nm, and 18001880 nm. The results are then weighted by the value of QUALITY and NDITH parameters, described in Sect. 1. By comparing the results from these two estimation methods we obtain an uncertainty in the spectral type of 1 subtype. Whenever the uncertainty is larger, a colon is added to the spectral subtype. A “p” is added to indicate a peculiar spectrum that fits to different subtypes over the wavelength range. This could be the case of non resolved binaries. An example of the resulting classification for a selection of the best quality spectra (few artifacts and high signal to noise, where possible) for each spectral subtype is plotted in Fig. 3 to show the spectral sequence. These spectra are plotted without any smoothing. The missing values are due to artifact removal (as shown in magenta in Fig. 1).
@@ 3.2 Spectra for Zhang’s compilation
We searched for Zhang’s compilation of UCD candidates in the Q1 released spectra using their object coordinates. Out of 4.3 million spectra, we found 458 objects matching within 1 arcsecond and 52 objects matching within 1 to 2.5 arcsecond of the coordinates. Together, these 510 objects account for 99% of Zhang’s compilation. An exhaustive comparison of these spectra with standard templates allowed us to confirm 224 UCDs. Table 2 shows the number of cross-matched objects, the number of confirmed dwarfs and the number of UCDs among them. In the case of EDF-N the confirmed UCDs represent 35% of the candidates with available spectra, given the limited depth of Zhang’s compilation in this field These numbers represent a minimum, as we excluded some candidates from confirmation owing to the poor SNR of their spectra. Future visits to the EDFs will allow us to increase the total number of confirmations. Fig. 4 shows the distribution of the number of objects per magnitude. The histograms are divided by field (North, South, and Fornax) to show the objects confirmed spectroscopically in each case. Note that from Zhang’s compilation, EDF-F and EDF-S (data from VHS) apparently cover fainter candidates than EDF-N (data from 2MASS). This is explained by the VHS -band magnitude, which is 3 magnitudes deeper than that from 2MASS -band. The UCDs with the best fits to the standard templates (negligible fit residuals) are selected as the ”UCD benchmarks”. There are 60 objects in this list, 37 late-M dwarfs, 21 L dwarfs, and 2 T dwarfs with magnitudes 14 19. Table 6 in Appendix B provides the full list. The rest of the confirmed objects include 49 late Ms, 90 Ls, and 6 Ts, most of them with large residuals and 12 that look like UCDs but cannot be reliably classified at the moment. The latter group of objects must wait for future visits to the EDFs which will allow a better classification, as the survey will gradually gain an additional depth of 2 magnitudes upon completion (Y. Euclid Collaboration: Mellier et al., 2024).
@@
Fig. 5 shows the spectra of E535151, the faintest L1 UCD from Zhang’s compilation found in Q1 data, together with that of E654087, the brightest L0 UCD from the same compilation. For this spectral type the span goes from 15.1 18.6.
3.3 Spectra retrieved by the spectral index search algorithm
@@
The 11 new late-L and T dwarfs discovered by the search algorithm were spectroscopically classified and their SpT is included in Table 1. There are UCDs for spectral types from L9 to T5. Three objects, E265716, E273062, and E271934 present a peculiar spectral energy distribution (signaled with a “p”), as shown in Fig. 6. E265716 could be a binary not revealed by the images, owing to its peculiar spectrum that fits a T4 in the -band and T2 in the -band. E271934 is a possible binary as the Euclid images show a close companion We are investigating this possibility by conducting follow-up observations (Muñoz Torres et al., in prep.). There are 5 additional T dwarfs found by the algorithm that were already known: one from Zhang’s compilation, two found by Žerjal’s et al. (in prep.) in the process of creating their photometric catalog, one from G. N. Mace et al. (2013), and one found as a QSO contaminant by Mohandasan et al. (in prep.). These objects are included in Table 7 (Appendix C) with the reference indicated next to the alias.